F(x) = square roots of (x+5) + 2, find the domain of f-1
Answer by Boreal(15235)

I am assuming this is the inverse function
y=sqrt(x+5)+2
x=sqrt(y+5)+2
sqrt(y+5)=x-2
y+5=x^2-4x+4
y=x^2-4x-1
The domain of the original function is x>=-5